OpenStudy (javsteratl):

WILL MEDAL.... Use substitution to answer the question below. 4x + 5y = 11 y = 3x ­ 13

OpenStudy (wkep1234):

alright you there?

OpenStudy (wkep1234):

see the 5y?

OpenStudy (wkep1234):

plug in y to that 5y: 4x+5(3x+13)=11

OpenStudy (wkep1234):

so then you will get 4x + 15x + 65 = 11

OpenStudy (wkep1234):

x = -54/19

OpenStudy (wkep1234):

or x = -2.84

OpenStudy (wkep1234):

medal?